OpenStudy (anonymous):

Suppose the population of a town is 1,100 and is growing 4% each year. a. Write an equation to model the population growth. b. Predict the population after 5 years.

OpenStudy (anonymous):

to increase a number by 4% multiply it by \(1.04\)

OpenStudy (anonymous):

so you should have \(1,100(1.04)^t\) as a formula. for 5 years replace t by 5

OpenStudy (anonymous):

\[\huge A=A_0e^{0.04t}\] \[\huge A=1100e^{0.04t}\]

OpenStudy (anonymous):

cyter formula for "continuous growth" but for 4% per year use the first one

OpenStudy (anonymous):

A.Y = 1,100x4^x;about 1,126,400 people B.y=4 x 1,100^x; about 22,000 people C.y=1,100 x 4^x; about 1,338 people D.y=1,100x1.04^x; about 1,338 people

OpenStudy (anonymous):

i think the last one looks a lot like what i wrote right?